# INSIGHT: Explainable Weakly-Supervised Medical Image Analysis

### v2

@@ 1 Introduction
Advances in medical imaging technology have enabled clinicians to extract critical insights from increasingly large and complex datasets. However, the size of medical images presents computational and analytical challenges: whole-slide images (WSIs) of histopathology slides can contain billions of pixels [3], and volumetric scans, such as CT or MRI, are composed of hundreds of slices. Processing such data end-to-end with deep neural networks is computationally infeasible. Instead, pipelines rely on aggregators, which synthesize local embeddings extracted from tiles (WSIs) or slices (volumes) into global predictions [5, 6, 23]. While this divide-and-conquer strategy is efficient, current methods often discard spatial information during feature aggregation and depend on post-hoc visualization tools, such as Grad-CAM [33], to generate interpretable heatmaps. These visualizations are prone to missing clinically significant features and introduce additional complexity.
@@
To address these limitations, we propose INSIGHT (Integrated Network for Segmentation and Interpretation with Generalized Heatmap Transmission), a novel weakly-supervised aggregator that embeds interpretability directly into its architecture. INSIGHT processes pre-trained patch- or slice-level embeddings through two complementary modules: a Detection Module, which captures fine-grained diagnostic features, and a Context Module, which suppresses irrelevant activations by incorporating broader spatial information. These outputs are fused to create interpretable internal heatmaps that naturally align with diagnostic regions. INSIGHT produces internal segmentation-quality heatmaps without requiring pixel- or voxel-level annotations as shown in Fig. LABEL:fig:visual-abstract.
@@
Weak supervision is essential in medical imaging, where the creation of detailed annotations is expensive and labor-intensive [3]. Fully supervised methods like U-Net [31] and DeepLab [7] rely on densely labeled data, which limits their scalability. In contrast, multiple instance learning (MIL) [4, 3] enables models to aggregate features from local regions using only image- or slide-level labels, avoiding the need for pixel-level annotations. This approach allows models to train on significantly larger datasets, as physicians typically diagnose patients without annotating individual image regions. For instance, the FDA-cleared Paige Prostate system demonstrates the clinical viability of weak supervision by achieving high diagnostic accuracy for prostate cancer detection [46, 3]. Similarly, MIL-based methods have effectively aggregated patch- or slice-level features to make predictions on WSIs and volumetric scans [5, 6].
@@
However, current weakly supervised methods face critical limitations. They require post-hoc visualization techniques, such as Grad-CAM [33] or class activation maps (CAM) [45], to produce interpretable saliency maps, adding complexity and often failing to localize small yet clinically essential details. These saliency maps can also be biased by dataset characteristics, such as organ co-occurrence, and often struggle to balance classification accuracy with spatial localization [10]. This leaves considerable room for improvement in designing aggregators that can better capture spatial dependencies and produce clinically meaningful outputs. INSIGHT addresses these challenges by directly embedding interpretability into its architecture, eliminating the reliance on post-hoc methods and bridging the gap between computational efficiency and clinical relevance.
@@
  • We demonstrate that INSIGHT achieves state-of-the-art performance on three challenging medical imaging datasets, excelling in both classification and segmentation tasks across CT and WSI modalities.
@@ 2.1 Weakly Supervised Medical Image Learning
Fully supervised methods such as U-Net [31] and DeepLab [7] have achieved high segmentation accuracy in medical imaging. However, their reliance on dense, pixel- or voxel-level annotations limits scalability in clinical workflows [3]. Weakly supervised learning mitigates this issue by enabling models to train on coarser annotations, such as image- or slide-level labels, which are more practical in large-scale clinical datasets.
Methods based on MIL have proven effective for WSIs and volumetric data, where detailed annotations are unavailable. MIL aggregates features from local patches or slices to predict global outcomes, bypassing the need for pixel-level labels. For example, the FDA-cleared Paige Prostate system demonstrated the clinical viability of MIL by achieving high diagnostic accuracy for prostate cancer detection using only slide-level labels [3, 46]. MIL-based methods have also been successfully applied to breast cancer subtyping from WSIs [5] and chemotherapy response prediction from volumetric CT scans [6].
While alternative weak labeling strategies, such as scribbles [19], bounding boxes [20], or point-level markers [16], reduce annotation demands, they still require explicit region annotations, which are rarely performed by physicians in clinical practice. INSIGHT builds on the MIL paradigm by utilizing only image-level labels for both classification and segmentation. Unlike existing methods, INSIGHT incorporates interpretability directly into its architecture, generating fine-grained heatmaps that align with diagnostic regions without requiring post-hoc visualization techniques.
@@ 2.2 Aggregation Techniques in Medical Imaging
Aggregation techniques are essential for efficiently processing large-scale medical imaging data, such as WSIs and volumetric scans, without requiring dense annotations. In WSIs, MIL-based methods treat each slide as a collection of patches and aggregate features for global predictions. For instance, MIL has been successfully applied to cancer detection using slide-level labels [3], while advancements like STAMP [25] and tissue-graph approaches [28] have improved WSI analysis by refining aggregation techniques to detect patterns and segment tissues.
In volumetric imaging, slice-level features are aggregated to form volume-level predictions. For example, multi-resolution systems like [32] use DenseVNet-generated features to improve classification of chest CT scans, while [43] developed an aggregation framework with label correction to enhance outcome predictions.
INSIGHT advances these techniques by preserving spatial resolution across slices or patches until the final pooling stage, enabling the generation of detailed, localized heatmaps. This addresses a key limitation of traditional aggregation methods, which often discard spatial information, reducing their ability to localize diagnostically relevant regions.
@@ 2.3 Explainable AI Methods in Medical Imaging
Explainable AI is crucial for ensuring model predictions in medical imaging are interpretable and clinically meaningful. Post-hoc methods such as class activation maps (CAM) [45] and Grad-CAM [33] are widely used to generate saliency maps that highlight regions associated with model predictions. For example, [42] developed high-resolution CAMs for thoracic CT scans, while [40] refined pseudo-labels for chest X-ray segmentation using CAMs.
Attention mechanisms have further advanced interpretability. Self-attention, as introduced in [39], has been combined with Grad-CAM to localize disease-specific regions [44], while dual-attention mechanisms, such as DA-CMIL [9], produce interpretable maps for detecting conditions like COVID-19 and bacterial pneumonia.
While these methods generate useful insights, they rely on post-hoc interpretability, which can misalign with the model’s decision-making process. INSIGHT eliminates this reliance by embedding interpretability directly into its architecture, producing calibrated, built-in heatmaps as part of its predictions. This ensures alignment between the model’s outputs and diagnostic reasoning.
@@ 2.4 Pre-training in Medical Image Analysis
Pre-trained models have significantly advanced medical imaging by providing robust feature extractors. ResNet-based architectures, such as those used in [11, 17], effectively capture local and intermediate features in weakly supervised frameworks. However, CNNs face limitations in capturing long-range dependencies, which are critical for analyzing complex medical images.
Vision Transformers (ViTs) [12] address these limitations by leveraging self-attention to capture broader spatial relationships. INSIGHT leverages pre-trained ViTs to extract tensor embeddings from WSI patches and radiology slices. By preserving spatial resolution, INSIGHT achieves robust performance across WSI and CT datasets, supporting fine-grained analysis and interpretability.
Figure 1: INSIGHT’s inputs and architecture. (a) Images are preprocessed to extract pre-trained features from each CT slice or WSI patch. (b) These features are processed through the Detection and Context modules to generate slice- or patch-level heatmaps by incorporating both fine-grained details and broader contextual information. (c) Heatmaps are aggregated across slices or patches to produce a binary prediction for each category along with interpretable heatmaps.
3 Method
Fig. 1 illustrates the architecture of INSIGHT, a weakly supervised framework for medical image analysis. INSIGHT processes pre-trained features from each slice in a volume or each patch in a WSI through dual modules—Detection and Context—to generate fine-grained heatmaps. These heatmaps are aggregated using a SmoothMax pooling strategy to produce categorical predictions, optimized using a combination of binary cross-entropy (BCE) and spectral decoupling losses. Below, we describe each component in detail.
@@ 3.1 Feature Extraction
To preserve spatial resolution and capture both local and global contextual information, INSIGHT uses pre-trained models specific to the input modality. For CT scan volumes, we use a ViT pre-trained with DINOv2 [26] on ImageNet. For pathology WSIs, we use UNI [8], a DINOv2 model pre-trained on over 100 million WSIs. The extracted 2D feature maps are denoted as $F\in\mathbb{R}^{c\times h\times w}$, where $c$, $h$, and $w$ are the feature depth, height, and width, respectively.
@@ 3.2 INSIGHT Architecture
Feature Adaptation. Pre-trained feature maps are transformed to domain-specific representations via a linear transformation layer:
$F_{t}=\mathbf{W}F+\mathbf{b},$ (1)
where $\mathbf{W}\in\mathbb{R}^{c^{\prime}\times c}$ and $\mathbf{b}\in\mathbb{R}^{c^{\prime}}$ adapt the features to the target dataset. The transformed features $F_{t}\in\mathbb{R}^{c^{\prime}\times h\times w}$ are then fed into two parallel modules: Detection and Context.
Detection Module. This module uses small-kernel convolutions to capture fine-grained details such as textures and edges, producing a heatmap $\mathcal{H}_{\texttt{Det}}$ that highlights potential regions of interest:
$\mathcal{H}_{\texttt{Det}}=\mathcal{M}_{\texttt{Det}}(F_{t}).$ (2)
Context Module. To reduce false positives from the Detection module, the Context module uses large-kernel convolutions to capture broader spatial context and suppress irrelevant activations. Its output, $\mathcal{H}_{\texttt{Con}}$, acts as a gating mechanism:
$\mathcal{H}_{\texttt{Con}}=\mathcal{M}_{\texttt{Con}}(F_{t}).$ (3)
Heatmap Fusion. The Detection and Context outputs are combined to generate a refined heatmap:
@@
where $\sigma$ is the sigmoid activation, and $\odot$ denotes element-wise multiplication. The resulting heatmap $\mathcal{H}$ assigns relevance scores to regions, aligning with the target labels.
3.3 Pooling and Classification
To produce a categorical outcome, heatmaps are aggregated across all slices or patches using SmoothMax pooling [22], a Boltzmann-weighted operator that emphasizes regions with strong activations:
$\hat{y}=\frac{\sum_{i}\mathcal{H}_{i}\cdot\exp(\alpha\cdot\mathcal{H}_{i})}{\sum_{i}\exp(\alpha\cdot\mathcal{H}_{i})},$ (5)
where $\hat{y}$ is the predicted likelihood of a positive outcome, $\mathcal{H}_{i}$ are heatmap activations, and $\alpha$ controls the weighting. For multi-label classification, the heatmap is extended to multiple channels, one per category, enabling simultaneous optimization for all labels.

### v3

@@ 1 Introduction
Advances in medical imaging technology have enabled clinicians to extract critical insights from increasingly large and complex datasets. However, the size of medical images presents computational and analytical challenges: whole-slide images (WSIs) of histopathology slides can contain billions of pixels (Campanella et al., 2019), and volumetric scans, such as CT or MRI, are composed of hundreds of slices. Processing such data end-to-end with deep neural networks is computationally infeasible. Instead, pipelines rely on aggregators, which synthesize local embeddings extracted from tiles (WSIs) or slices (volumes) into global predictions (Casson et al., 2024; Chang et al., 2022; Mahmood et al., 2021). While this divide-and-conquer strategy is efficient, current methods often discard spatial information during feature aggregation and depend on post-hoc visualization tools, such as Grad-CAM (Selvaraju et al., 2019), to generate interpretable heatmaps. These visualizations are prone to missing clinically significant features and introduce additional complexity.
@@
To address these limitations, we propose INSIGHT (Integrated Network for Segmentation and Interpretation with Generalized Heatmap Transmission), a novel weakly-supervised aggregator that embeds interpretability directly into its architecture. INSIGHT processes pre-trained patch- or slice-level embeddings through two complementary modules: a Detection Module, which captures fine-grained diagnostic features, and a Context Module, which suppresses irrelevant activations by incorporating broader spatial information. These outputs are fused to create interpretable internal heatmaps that naturally align with diagnostic regions. INSIGHT produces internal segmentation-quality heatmaps without requiring pixel- or voxel-level annotations as shown in Fig. 1.
Figure 1: Qualitative visualization of interpretable heatmaps generated by INSIGHT. (a) Original whole-slide image (WSI) with annotated tumor regions outlined in green. (b) WSI-level heatmap produced by our method, highlighting predicted tumor regions. (c) Zoomed-in view of a representative lesion with expert annotation. (d) Corresponding zoomed-in heatmap showing precise localization of the lesion by INSIGHT. All heatmaps are generated using only image-level categorical labels.
@@
Weak supervision is essential in medical imaging, where the creation of detailed annotations is expensive and labor-intensive (Campanella et al., 2019). Fully supervised methods like U-Net (Ronneberger et al., 2015) and DeepLab (Chen et al., 2017) rely on densely labeled data, which limits their scalability. In contrast, multiple instance learning (MIL) (Carbonneau et al., 2018; Campanella et al., 2019) enables models to aggregate features from local regions using only image- or slide-level labels, avoiding the need for pixel-level annotations. This approach allows models to train on significantly larger datasets, as physicians typically diagnose patients without annotating individual image regions. For instance, the FDA-cleared Paige Prostate system demonstrates the clinical viability of weak supervision by achieving high diagnostic accuracy for prostate cancer detection (Zhu et al., 2024; Campanella et al., 2019). Similarly, MIL-based methods have effectively aggregated patch- or slice-level features to make predictions on WSIs and volumetric scans (Casson et al., 2024; Chang et al., 2022).
@@
However, current weakly supervised methods face critical limitations. They require post-hoc visualization techniques, such as Grad-CAM (Selvaraju et al., 2019) or class activation maps (CAM) (Zhou et al., 2015), to produce interpretable saliency maps, adding complexity and often failing to localize small yet clinically essential details. These saliency maps can also be biased by dataset characteristics, such as organ co-occurrence, and often struggle to balance classification accuracy with spatial localization (Cohen et al., 2020). This leaves considerable room for improvement in designing aggregators that can better capture spatial dependencies and produce clinically meaningful outputs. INSIGHT addresses these challenges by directly embedding interpretability into its architecture, eliminating the reliance on post-hoc methods and bridging the gap between computational efficiency and clinical relevance.
@@
  • We demonstrate that INSIGHT achieves state-of-the-art performance on three challenging medical imaging datasets, excelling in both classification and segmentation tasks across CT and WSI modalities using only image-level categorical labels.
@@
Generalizable Insights about Machine Learning in the Context of Healthcare
INSIGHT demonstrates that fine-grained interpretability and strong predictive performance are not mutually exclusive in weakly supervised settings. By maintaining spatial resolution throughout the feature processing pipeline and embedding interpretability directly into the architecture, our method provides a general design principle for scalable and explainable medical AI systems. This architectural approach—preserving locality while incorporating contextual suppression—can be extended beyond pathology and CT to other modalities such as MRI or ultrasound. More broadly, INSIGHT challenges the common reliance on post-hoc attention visualization, offering a viable alternative that integrates decision explanation into the prediction mechanism itself.
For the clinical radiologist or pathologist, INSIGHT produces built-in heatmaps that align with diagnostic regions without requiring pixel-level annotations. These interpretable outputs help clinicians verify model predictions and identify overlooked regions, while traditional pixel-level annotation remains both labor-intensive and time-consuming. By reducing reliance on such dense annotations and producing visual explanations, INSIGHT can accelerate image review and enhance diagnostic confidence, especially for subtle or ambiguous cases.
Figure 2: Semantic segmentation for medical data often requires expensive annotations that are not done routinely in clinical practice, whereas INSIGHT only needs Categorical labels (left). Unlike other weakly-labeled methods, INSIGHT has per-label heatmaps built into the architecture as a form of inductive bias, enabling it to directly produce heatmaps that closely align with ground-truth diagnostic regions without the use of post-hoc methods like Grad-CAM (right).
@@ 2.1 Weakly Supervised Medical Image Learning
Fully supervised methods such as U-Net (Ronneberger et al., 2015) and DeepLab (Chen et al., 2017) have achieved high segmentation accuracy in medical imaging. However, their reliance on dense, pixel- or voxel-level annotations limits scalability in clinical workflows (Campanella et al., 2019). Weakly supervised learning mitigates this issue by enabling models to train on coarser annotations, such as image- or slide-level labels, which are more practical in large-scale clinical datasets. Methods based on MIL have proven effective for WSIs and volumetric data, where detailed annotations are unavailable. MIL aggregates features from local patches or slices to predict global outcomes, bypassing the need for pixel-level labels. For example, the FDA-cleared Paige Prostate system demonstrated the clinical viability of MIL by achieving high diagnostic accuracy for prostate cancer detection using only slide-level labels (Campanella et al., 2019; Zhu et al., 2024). MIL-based methods have also been successfully applied to breast cancer subtyping from WSIs (Casson et al., 2024) and chemotherapy response prediction from volumetric CT scans (Chang et al., 2022).
While alternative weak labeling strategies, such as scribbles (Liu et al., 2022), bounding boxes (Lu et al., 2023), or point-level markers (Laradji et al., 2020), reduce annotation demands, they still require explicit region annotations, which are rarely performed by physicians in clinical practice. INSIGHT builds on the MIL paradigm by utilizing only image-level labels for both classification and segmentation. Unlike existing methods, INSIGHT incorporates interpretability directly into its architecture, generating fine-grained heatmaps that align with diagnostic regions without requiring post-hoc visualization techniques.
@@ 2.2 Aggregation Techniques in Medical Imaging
Aggregation techniques are essential for efficiently processing large-scale medical imaging data, such as WSIs and volumetric scans, without requiring dense annotations. In WSIs, MIL-based methods treat each slide as a collection of patches and aggregate features for global predictions. For instance, MIL has been successfully applied to cancer detection using slide-level labels (Campanella et al., 2019), while advancements like STAMP (Nahhas et al., 2023) and tissue-graph approaches (Pati et al., 2023) have improved WSI analysis by refining aggregation techniques to detect patterns and segment tissues. In volumetric imaging, slice-level features are aggregated to form volume-level predictions. For example, multi-resolution systems like Saha et al. (2020) use DenseVNet-generated features to improve classification of chest CT scans, while Ye et al. (2022) developed an aggregation framework with label correction to enhance outcome predictions.
INSIGHT advances these techniques by preserving spatial resolution across slices or patches until the final pooling stage, enabling the generation of detailed, localized heatmaps. This addresses a key limitation of traditional aggregation methods, which often discard spatial information, reducing their ability to localize diagnostically relevant regions. Existing methods aim at enhancing contextual understanding in feature aggregation, but they each have inherent limitations. For instance, FPN (Lin et al., 2017) is designed for hierarchical multi-scale feature fusion but lacks an explicit mechanism to independently optimize local and global features, which is crucial for WSI. On the other hand, Dual Attention Network (Fu et al., 2019) relies on self-attention mechanisms to model long-range dependencies across spatial and channel dimensions, leading to quadratic complexity and computationally expensive for high-resolution medical images such as WSI. In contrast, unlike FPN which depends on predefined hierarchical feature aggregation, INSIGHT’s Context module efficiently refines activations within the same feature scale. It also allows separate optimizations on local and global representations. Meanwhile, compared to Dual Attention Nets, INSIGHT avoids the quadratic complexity of self-attention with larger convolution kernel size, making it significantly more efficient for processing whole-slide images.
@@ 2.3 Explainable AI Methods in Medical Imaging
Explainable AI is crucial for ensuring model predictions in medical imaging are interpretable and clinically meaningful. Post-hoc methods such as class activation maps (CAM) (Zhou et al., 2015) and Grad-CAM (Selvaraju et al., 2019) are widely used to generate saliency maps that highlight regions associated with model predictions. For example, Xie et al. (2023) developed high-resolution CAMs for thoracic CT scans, and Viniavskyi et al. (2020) refined pseudo-labels for chest X-ray segmentation using CAMs. Attention mechanisms have further advanced interpretability. Self-attention, as introduced in (Vaswani et al., 2023), has been combined with Grad-CAM to localize disease-specific regions (Zhang et al., 2022), while dual-attention mechanisms, such as DA-CMIL (Chikontwe et al., 2021), produce interpretable maps for detecting conditions like COVID-19 and bacterial pneumonia.
While these methods generate useful insights, they rely on post-hoc interpretability, which can misalign with the model’s decision-making process. Few existing methods directly integrate interpretability into their model architectures. One example is Shallow-ProtoPNet (Singh et al., 2025), a prototype-based network designed for fully transparent inference on small-scale datasets such as chest X-rays. In contrast, INSIGHT is built to scale to large, high-resolution inputs like CT volumes and WSIs, addressing the computational and spatial challenges that arise in real-world clinical imaging. INSIGHT eliminates this reliance by embedding interpretability directly into its architecture, producing calibrated, built-in heatmaps as part of its predictions, as shown in Fig. 3. ensures alignment between the model’s outputs and diagnostic reasoning.
@@ 2.4 Pre-training in Medical Image Analysis
Pre-trained models have significantly advanced medical imaging by providing robust feature extractors. ResNet-based architectures, such as those used in (Courtiol et al., 2020; Lin et al., 2022), effectively capture local and intermediate features in weakly supervised frameworks.
On the other hand, Vision Transformers (ViTs) (Dosovitskiy et al., 2021) leverage self-attention to capture broader spatial relationships. Self-supervised foundational models are typically pre-trained on massive unlabeled image datasets, enabling the learning of meaningful representations without manual annotations. They produce high-quality features that are highly compatible with downstream tasks including those based on aggregation.
For example, UNI (Chen et al., 2024) is trained on over $100$ M WSIs using a DINOv2-style (Oquab et al., 2024) self-supervised learning framework and provides pathology-specific representations optimized for fine-grained localization. Virchow 2 (Zimmermann et al., 2024), a $632$-million-parameter ViT-H model, is pre-trained on $3.1$ M WSIs with domain-specific augmentations tailored for histopathology tasks. INSIGHT leverages pre-trained ViTs to extract tensor embeddings from WSI patches and radiology slices. By preserving spatial resolution, INSIGHT achieves robust performance across WSI and CT datasets, supporting fine-grained analysis and interpretability.
Figure 3: Overview of the INSIGHT framework. (a) Input images (WSIs or CT volumes) are preprocessed and transformed into spatial embeddings using a pretrained encoder. (b) Each slice or patch embedding is processed by INSIGHT, which consists of a detection module to capture fine-grained signals and a context suppression module to reduce false positives. This produces patch-level heatmaps for each category, which are then aggregated to generate whole-slide or volume-level heatmaps. Final predictions for each category are obtained via SmoothMax pooling over the heatmaps. Throughout this process, spatial resolution is preserved, encouraging the model to produce reliable and interpretable heatmaps.
3 Methods
Fig. 3 illustrates the high-level overview architecture of INSIGHT. INSIGHT processes pre-trained features from each slice in a volume or each patch in a WSI through dual modules—Detection and Context—to generate fine-grained heatmaps. These heatmaps are aggregated using a SmoothMax pooling strategy to produce categorical predictions, optimized using a combination of binary cross-entropy (BCE) and spectral decoupling losses. Below, we describe each component in detail.
@@ 3.1 Feature Extraction
To preserve spatial resolution and capture both local and global contextual information, INSIGHT uses pre-trained models specific to the input modality, as illustrated in Fig 3(a). To demonstrate that INSIGHT is compatible with a variety of foundation models across modalities, we evaluate it with multiple state-of-the-art encoders. For CT scan volumes, we use a ViT pre-trained with DINOv2 (Oquab et al., 2024) on ImageNet. For pathology WSIs, we evaluate INSIGHT with both UNI (Chen et al., 2024) and Virchow 2 (Zimmermann et al., 2024). UNI is pre-trained using DINOv2-style self-supervision on over 100 million WSIs, enabling it to capture rich visual features tailored to histopathology. Virchow 2 is a large ViT-H model with 632 million parameters, trained on 3.1 million WSIs using domain-specific augmentations optimized for pathology. In all cases, the extracted feature maps are denoted as $F\in\mathbb{R}^{c\times h\times w}$, where $c$, $h$, and $w$ are the feature depth, height, and width, respectively.
@@ 3.2 INSIGHT Architecture
Heatmap Fusion. The core of INSIGHT lies in its architectural design which introduces an inductive bias that integrates both local and global spatial features. As illustrated in Fig. 3(b), each input embedding is processed by two parallel modules: the Detection module and the Context modules, which are combined to generate a refined heatmap:
@@
where $\sigma$ is the sigmoid activation, and $\odot$ denotes element-wise multiplication. The resulting heatmap $\mathcal{H}$ assigns relevance scores to regions, aligning with the target labels. Detection module uses small-kernel convolutions to capture fine-grained details such as textures and edges, producing a heatmap $\mathcal{H}_{\texttt{Det}}$ that highlights potential regions of interest. To mitigate false positives, the Context module uses larger kernels to capture broader spatial context and suppress irrelevant activations. To better align the pre-trained features with the downstream task, we apply a linear projection before passing the embeddings to both modules.
To isolate high-saliency regions within the fused heatmap, we apply dynamic thresholding with Otsu’s algorithm (Otsu, 1979):
$\mathcal{H^{\prime}}=\mathcal{H}\cdot\mathbb{I}(\mathcal{H}>T),$ (2)
where $T=\arg\min_{t}\sigma_{w}^{2}(t)$ minimizes intra-class variance and $\mathbb{I}$ is the indicator function. This step enhances localization by filtering out low-activation background regions.
Pooling and Classification. To produce a categorical outcome, heatmaps are aggregated across all slices or patches using SmoothMax pooling (Maddison et al., 2016), which is a Boltzmann-weighted operator that emphasizes regions with strong activations:
$\hat{y}=\frac{\sum_{i}\mathcal{H^{\prime}}_{i}\cdot\exp(\alpha\cdot\mathcal{H^{\prime}}_{i})}{\sum_{i}\exp(\alpha\cdot\mathcal{H^{\prime}}_{i})},$ (3)
where $\hat{y}$ is the predicted likelihood of a positive outcome, $\mathcal{H^{\prime}}_{i}$ are heatmap activations, and $\alpha$ controls the sharpness of the weighting. For multi-label classification, the heatmap is extended across channels, with each channel representing a separate category, allowing joint optimization over multiple labels.
We empirically tested several other pooling methods, including max pooling and LP pooling to compare their effectiveness. Max pooling considers only the strongest activation, which limits spatial gradient flow during training and leads to suboptimal heatmap quality. LP pooling ($p\geq 2$)incorporates more values but often produces outputs outside the $[0,1]$ range, requiring post-processing (e.g., output clipping or sparse regularization (Zhou et al., 2021)), which yielded no empirical gains. In contrast, SmoothMax pooling provides a principled trade-off between local emphasis and global context through exponential weighting.